### v1

# Uncertainty, von Neumann entropy and squeezing in bipartite state of two-level atoms

Abstract
[…]
pacs:
03.67.Mn, 03.65.Ud, 42.50.Dv, 42.50.-p
I I INTRODUCTION
Uncertainties in the simultaneous measurements of incompatible observables of quantum systems make quantum systems drastically different from classical ones. Uncertainty principle, first formulated by Heisenberg Heisenberg forms a central part of our understanding of quantum mechanics. Many works have been done on further developments and applications of uncertainty principle with the advent and progress of quantum information theory Robertson -Horodecki1 .
Von Neumann entropy, also called entanglement entropy, plays a crucial role in the quantification of quantum entanglement. Entropic manifestation of quantum entanglement forms a significant part of quantum information theory Horodecki1 .
Spin squeezing is an important experimental tool that have applications in the detection of quantum entanglement, high-precision metrology, atomic clocks, quantum Fisher information, quantum chaos, quantum phase transitions, Bose-Einstein condensates, and ion-trap experiments in quantum optics Jian .
In this paper we consider a general entangled state of two two-level atoms. We show that though the von Neumann entropies of the entangled subsystems are always greater than that of the system as a whole, thus violating the entropic inequalities Horodecki1 , but the uncertainties of the subsystems are less than the system as a whole for certain range of the superposing constants of the quantum state. This is in contradiction with the consensus that more is the entropy, more is the uncertainty and more is the disorder. We also establish analytically a general relationship between the von Neumann entropies of the subsystems and the uncertainties, spin squeezing and spectroscopic squeezing parameters of the system as a whole. Now, since the spin squeezing and spectroscopic squeezing are experimentally measurable, so, this provides an operational measure of the entanglement entropy of the subsystems of two-level atoms.
In section II we briefly discuss the two-level atoms and pseudo-spin operators. In section III we calculate the uncertainties of the pseudo-spin operators in two mutually orthogonal directions in a plane perpendicular to the mean spin vector of the general entangled state of two two-level atoms, both for the composite state and the subsystems. We show that though the von Neumann entropy of the composite state is less than those of the subsystems, but, the uncertainties of the composite state are greater than those of the subsystems for certain range of the superposing constants of the quantum state. In section IV we establish general relationships between the von Neumann entropies of the subsystems and the uncertainties, spin squeezing and spectroscopic squeezing parameters of the composite state. In section V we present the conclusion.
I.1 II TWO-LEVEL ATOM, PSEUDO-SPIN OPERATORS
An atom has many electronic energy levels, but when it interacts with an external monochromatic electromagnetic radiation, the atom makes a transition from one of its energy level to the other. In this case, we mainly concentrate on those two energy levels and hence the atom is called as a two-level atom. Now, if among the assembly of $N$ such two-level atoms, the $n$-th atom has the upper and lower energy levels, denoted as $|u_{n}\rangle$ and $|l_{n}\rangle$, respectively, then, we can construct the pseudo-spin operators (with $\hbar=1$),
$\hat{J}_{n_{x}}$ $=$ $(1/2)\big{(}|u_{n}\rangle\langle l_{n}|+|l_{n}\rangle\langle u_{n}|\big{)},$ (1)
$\hat{J}_{n_{y}}$ $=$ $(-i/2)\big{(}|u_{n}\rangle\langle l_{n}|-|l_{n}\rangle\langle u_{n}|\big{)},$ (2)
$\hat{J}_{n_{z}}$ $=$ $(1/2)\big{(}|u_{n}\rangle\langle u_{n}|-|l_{n}\rangle\langle l_{n}|\big{)},$ (3)
such that,
$[\hat{J}_{n_{x}},\hat{J}_{n_{y}}]=i\hat{J}_{n_{z}},$ (4)
and two more relations with cyclic changes in $x$, $y$ and $z$ Itano . For the entire system of two two-level atoms, we have collective pseudo-spin operators,
$\hat{J}_{x}=\sum_{i=1}^{2}\hat{J}_{i_{x}},\hat{J}_{y}=\sum_{i=1}^{2}\hat{J}_{i_{y}},\hat{J}_{z}=\sum_{i=1}^{2}\hat{J}_{i_{z}},$ (5)
where each term in the above summations is multiplied by the identity operators for all of the other atoms Itano , Sakurai .
The individual atomic operators satisfy
$\big{[}\hat{J}_{1_{x}},\hat{J}_{2_{y}}\big{]}$ $=$ $0,~{}\big{[}\hat{J}_{1_{x}},\hat{J}_{1_{y}}\big{]}=i\hat{J}_{1_{z}},$
$\big{[}\hat{J}_{2_{x}},\hat{J}_{2_{y}}\big{]}$ $=$ $i\hat{J}_{2_{z}},...$ (6)
where the subscripts $1$ and $2$ are for atoms $1$ and $2$ respectively.
As a consequence of these commutation relations, the collective pseudo-spin operators $\hat{J}_{x}$, $\hat{J}_{y}$ and $\hat{J}_{z}$ satisfy,
$[\hat{J}_{x},\hat{J}_{y}]=i\hat{J}_{z}$ (7)
and two more relations with cyclic changes in $x$, $y$ and $z$.
The eigenvectors of $\hat{J}_{n_{z}}$ are denoted as $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$ with eigenvalues $\frac{1}{2}$ and $-\frac{1}{2}$ respectively.
The use of these collective spin operators as discussed in Eqs. (5) are widely used in the problems of quantum optics Korbicz1 -Miyake2 .
I.2 III UNCERTAINTIES AND VON NEUMANN ENTROPY IN A SYSTEM OF TWO TWO-LEVEL ATOMS
We, now, consider a system of two entangled two-level atoms, say atoms $a$ and $b$. A general pure state of this system in Schmidt decomposition is given as
$|\psi_{ab}\rangle=\sum_{i=1}^{2}c_{i}|e^{i}_{a}\rangle\otimes|e^{i}_{b}\rangle,$ (8)
where $|\psi_{ab}\rangle$in $H_{a}\otimes H_{b}$, $c_{i}$ are the Schmidt coefficients and $\{|e^{i}_{a}\rangle\otimes|e^{i}_{b}\rangle\}$ is a orthonormal product basis. The Schmidt coefficients $c_{i}$ satisfy
$\sum_{i=1}^{2}{c_{i}}^{2}=1.$ (9)
For the system of two two-level atoms we take the orthonormal product basis, made up of generalized superpositions of $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$, as
$\Big{\{}\big{(}c_{3}|\frac{1}{2}\rangle+c_{4}|-\frac{1}{2}\rangle\big{)}\otimes\big{(}c_{5}|\frac{1}{2}\rangle+c_{6}|-\frac{1}{2}\rangle\big{)}\Big{\}}$
and
$\Big{\{}\big{(}c_{4}|\frac{1}{2}\rangle-c_{3}|-\frac{1}{2}\rangle\big{)}\otimes\big{(}c_{6}|\frac{1}{2}\rangle-c_{5}|-\frac{1}{2}\rangle\big{)}\Big{\}},$
where $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$ are the eigenvectors of $\hat{J}_{a_{z}}$ and $\hat{J}_{b_{z}}$ with eigenvalues $\frac{1}{2}$ and $-\frac{1}{2}$ respectively and for the sake of simplicity $c_{3}$, $c_{4}$, $c_{5}$ and $c_{6}$ are assumed to be real. Since the basis vectors are normalized, we have
${c_{3}}^{2}+{c_{4}}^{2}$ $=$ $1,$ (10)
${c_{5}}^{2}+{c_{6}}^{2}$ $=$ $1$ (11)
Therefore, the complete form of $|\psi_{ab}\rangle$ is
$|\psi_{ab}\rangle=c_{1}\bigg{[}\big{(}c_{3}\Big{|}\frac{1}{2}\Big{\rangle}+c_{4}\Big{|}-\frac{1}{2}\Big{\rangle}\big{)}\otimes$
$\big{(}c_{5}\Big{|}\frac{1}{2}\Big{\rangle}+c_{6}\Big{|}-\frac{1}{2}\Big{\rangle}\big{)}\bigg{]}+$
$c_{2}\bigg{[}\big{(}c_{4}\Big{|}\frac{1}{2}\Big{\rangle}-c_{3}\Big{|}-\frac{1}{2}\Big{\rangle}\big{)}$
$\otimes\big{(}c_{6}\Big{|}\frac{1}{2}\Big{\rangle}-c_{5}\Big{|}-\frac{1}{2}\Big{\rangle}\big{)}\bigg{]}$ (12)
The above quantum state can be written as
$|\psi_{ab}\rangle=\Big{(}c_{1}c_{3}c_{5}+c_{2}c_{4}c_{6}\Big{)}\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+$
$\Big{(}c_{1}c_{3}c_{6}-c_{2}c_{4}c_{5}\Big{)}\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}+$
$\Big{(}c_{1}c_{4}c_{5}-c_{2}c_{3}c_{6}\Big{)}\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+$
$\Big{(}c_{1}c_{4}c_{6}+c_{2}c_{3}c_{5}\Big{)}\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}$ (13)
For making the calculations simple we let
$\Big{(}c_{1}c_{3}c_{5}+c_{2}c_{4}c_{6}\Big{)}$ $=$ $A,$ (14)
$\Big{(}c_{1}c_{3}c_{6}-c_{2}c_{4}c_{5}\Big{)}$ $=$ $B$ (15)
$\Big{(}c_{1}c_{4}c_{5}-c_{2}c_{3}c_{6}\Big{)}$ $=$ $C,$ (16)
$\Big{(}c_{1}c_{4}c_{6}+c_{2}c_{3}c_{5}\Big{)}$ $=$ $D,$ (17)
where
$A^{2}+B^{2}+C^{2}+D^{2}=1.$ (18)
In order to study the spin squeezing and spectroscopic squeezing properties of the system we need to calculate the coordinate frame independent inherent uncertainties in the components of $\hat{\mathbf{J}}$ in two mutually orthogonal directions, in a plane perpendicular to the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ of the system Kitagawa ; Wineland2 , where
$\langle\hat{\mathbf{J}}\rangle=\langle\hat{J}_{n_{1}}\rangle\hat{n}_{1}+\langle\hat{J}_{n_{2}}\rangle\hat{n}_{2}+\langle\hat{J}_{n_{3}}\rangle\hat{n}_{3}.$ (19)
$\hat{n}_{1}$, $\hat{n}_{2}$ and $\hat{n}_{3}$ are three mutually orthogonal unit vectors. The expectation values as mentioned above are to be calculated over the quantum state in Eq. (12).
Now, to understand it concretely we take the directions of $\hat{n}_{1}$, $\hat{n}_{2}$ and $\hat{n}_{3}$ along the positive $x$, $y$ and $z$ axes respectively so that $\langle\hat{\mathbf{J}}\rangle$ is
$\langle\hat{\mathbf{J}}\rangle=\langle\hat{J}_{x}\rangle\hat{i}+\langle\hat{J}_{y}\rangle\hat{j}+\langle\hat{J}_{z}\rangle\hat{k},$ (20)
and hence the uncertainties are the standard deviations
$\Delta J_{x,y}=\sqrt{\langle\psi_{ab}|{\hat{J}_{x,y}}^{2}|\psi_{ab}\rangle-\langle\psi_{ab}|\hat{J}_{x,y}|\psi_{ab}\rangle^{2}}.$ (21)
Now, since for an arbitrary quantum state, as that given in Eq. (12) the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ points in an arbitrary direction in space, we conventionally rotate the coordinate system $\{x,y,z\}$ to $\{x^{\prime},y^{\prime},z^{\prime}\}$, such that $\langle\hat{\bf{J}}\rangle$ points along the $z^{\prime}$ axis and calculate the standard deviations of $\hat{J}_{x^{\prime}}$ and $\hat{J}_{y^{\prime}}$ for the state $|\psi_{ab}\rangle$. These are
$\Delta J_{x^{\prime},y^{\prime}}=\sqrt{\langle\psi_{ab}|{\hat{J}_{x^{\prime},y^{\prime}}}^{2}|\psi_{ab}\rangle-\langle\psi_{ab}|\hat{J}_{x^{\prime},y^{\prime}}|\psi_{ab}\rangle^{2}}.$ (22)
Now, since the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ after rotation of the coordinate frame points along the $z^{\prime}$ axis, we have
$\langle\psi_{ab}|\hat{J}_{x^{\prime}}|\psi_{ab}\rangle=\langle\psi_{ab}|\hat{J}_{y^{\prime}}|\psi_{ab}\rangle=0,$ (23)
and hence, the standard deviations reduce to
$\Delta J_{x^{\prime},y^{\prime}}=\sqrt{\langle\psi_{ab}|{\hat{J}_{x^{\prime},y^{\prime}}}^{2}|\psi_{ab}\rangle}.$ (24)
We now proceed to calculate $\Delta J_{x^{\prime},y^{\prime}}$. We find that
$\langle\psi_{ab}|\hat{J}_{x}|\psi_{ab}\rangle$ $=$ $(A+D)(B+C)$ (25)
$\langle\psi_{ab}|\hat{J}_{y}|\psi_{ab}\rangle$ $=$ $0,$ (26)
$\langle\psi_{ab}|\hat{J}_{z}|\psi_{ab}\rangle$ $=$ $A^{2}-D^{2}.$ (27)
Therefore, the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ lies in the $x-z$ plane. To align it along the $z^{\prime}$ direction we perform the following rotation.
$\hat{J}_{x^{\prime}}$ $=$ $\hat{J}_{x}\cos\theta-\hat{J}_{z}\sin\theta$ (28)
$\hat{J}_{y^{\prime}}$ $=$ $\hat{J}_{y}$ (29)
$\hat{J}_{z^{\prime}}$ $=$ $\hat{J}_{x}\sin\theta+\hat{J}_{z}\cos\theta,$ (30)
where,
$\cos\theta$ $=$ $\frac{\langle\hat{J}_{z}\rangle}{|\langle\hat{\mathbf{J}}\rangle|}.$ (31)
It can be verified using Eqs. (26), (28), (29) and (31) that
$\langle\psi_{ab}|\hat{J}_{x^{\prime}}|\psi_{ab}\rangle$ $=$ 00 0 (32)
$\langle\psi_{ab}|\hat{J}_{y^{\prime}}|\psi_{ab}\rangle$ $=$ $0,$ (33)
and hence the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ is now along the $z^{\prime}$-axis.
Now, using Eqs. (28) and (31), we get
$\langle\hat{J}_{x^{\prime}}^{2}\rangle=\frac{1}{|\langle\hat{\mathbf{J}}\rangle|^{2}}\Big{[}\langle\hat{J}_{x}^{2}\rangle\langle\hat{J}_{z}\rangle^{2}+\langle\hat{J}_{z}^{2}\rangle\langle\hat{J}_{x}\rangle^{2}$
$-\langle\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}\rangle\langle\hat{J}_{x}\rangle\langle\hat{J}_{z}\rangle\Big{]}.$ (34)
We find that for the quantum state in Eq. (12),
$\langle\psi_{ab}|\hat{J}_{x}^{2}|\psi_{ab}\rangle=\frac{1}{2}\Big{(}1+2AD+2BC\Big{)},$ (35)
$\langle\psi_{ab}|\hat{J}_{z}^{2}|\psi_{ab}\rangle=A^{2}+D^{2},$ (36)
$\langle\psi_{ab}|\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}|\psi_{ab}\rangle=(A-D)(B+C).$
Therefore, using Eqs. (25), (32), (34) to (I.2), we get
$\langle\psi_{ab}|\hat{J}_{x^{\prime}}^{2}|\psi_{ab}\rangle=\frac{1}{2}\big{(}A+D\big{)}^{2}$ (38)
So, using Eqs. (24) and (38), we get
$\Delta{J}_{x^{\prime}}^{2}=\frac{1}{2}\big{(}A+D\big{)}^{2}.$ (39)
Similarly,
$\langle\psi_{ab}|\hat{J}_{y}^{2}|\psi_{ab}\rangle=\frac{1}{2}\Big{(}1-2AD+2BC\Big{)}.$ (40)
Therefore, using Eqs. (24), (26), (29), (33) and (40), we get
$\Delta{J}_{y^{\prime}}^{2}$ $=$ $\langle\psi_{ab}|\hat{J}_{y^{\prime}}^{2}|\psi_{ab}\rangle$ (41)
$=$ $\frac{1}{2}\Big{(}1-2AD+2BC\Big{)}.$
Now, using Eqs. (14), (15), (16), (17), (39) and (41), we get
$\Delta{J}_{x^{\prime}}=\sqrt{\Big{(}\frac{1}{2}+c_{1}c_{2}\Big{)}}\Big{(}c_{3}c_{5}+c_{4}c_{6}\Big{)},$ (42)
$\Delta{J}_{y^{\prime}}=\sqrt{\Big{(}\frac{1}{2}-c_{1}c_{2}\Big{)}}.$ (43)
These are the coordinate frame independent inherent uncertainties of the system.
We now proceed to calculate similar uncertainties in the atomic operators $\hat{J}_{a_{x^{\prime}}}$, $\hat{J}_{a_{y^{\prime}}}$, $\hat{J}_{b_{x^{\prime}}}$ and $\hat{J}_{b_{y^{\prime}}}$ of atoms $a$ and $b$ respectively.
Now the reduced density matrix $\hat{\rho}_{a}$ for atom $a$ is
$\hat{\rho}_{a}=c_{1}^{2}\Big{[}c_{3}\Big{|}\frac{1}{2}\Big{\rangle}+c_{4}\Big{|}-\frac{1}{2}\Big{\rangle}\Big{]}\Big{[}\Big{\langle}\frac{1}{2}\Big{|}c_{3}+$
$\Big{\langle}-\frac{1}{2}\Big{|}c_{4}\Big{]}+c_{2}^{2}\Big{[}c_{4}\Big{|}\frac{1}{2}\Big{\rangle}-c_{3}\Big{|}-\frac{1}{2}\Big{\rangle}\Big{]}$
$\Big{[}\Big{\langle}\frac{1}{2}\Big{|}c_{4}-\Big{\langle}-\frac{1}{2}\Big{|}c_{3}\Big{]}$ (44)
We find that
$\langle\hat{J}_{a_{x}}\rangle=tr(\hat{\rho}_{a}\hat{J}_{a_{x}})=c_{3}c_{4}(c_{1}^{2}-c_{2}^{2}),$ (45)
$\langle\hat{J}_{a_{y}}\rangle=tr(\hat{\rho}_{a}\hat{J}_{a_{y}})=0,$ (46)
$\langle\hat{J}_{a_{z}}\rangle=tr(\hat{\rho}_{a}\hat{J}_{a_{z}})=\frac{1}{2}(c_{1}^{2}-c_{2}^{2})(c_{3}^{2}-c_{4}^{2}).$
So, the mean spin vector $\langle\hat{\mathbf{J}}_{a}\rangle$ lies in the $x-z$ plane. Therefore, we perform the rotation
$\hat{J}_{a_{x^{\prime}}}$ $=$ $\hat{J}_{a_{x}}\cos\theta_{1}-\hat{J}_{a_{z}}\sin\theta_{1}$ (48)
$\hat{J}_{a_{y^{\prime}}}$ $=$ $\hat{J}_{a_{y}}$ (49)
$\hat{J}_{a_{z^{\prime}}}$ $=$ $\hat{J}_{a_{x}}\sin\theta_{1}+\hat{J}_{a_{z}}\cos\theta_{1},$ (50)
where
$\cos\theta_{1}$ $=$ $\frac{\langle\hat{J}_{a_{z}}\rangle}{|\langle\hat{\mathbf{J}}_{a}\rangle|}.$ (51)
It can be checked that
$\langle\hat{J}_{a_{x^{\prime}}}\rangle$ $=$ 00 0 (52)
$\langle\hat{J}_{a_{y^{\prime}}}\rangle$ $=$ $0,$ (53)
and hence the mean spin vector $\langle\hat{\mathbf{J}}_{a}\rangle$ is now along the $z^{\prime}$-axis. Therefore,
$\Delta J_{a_{x^{\prime},y^{\prime}}}=\sqrt{\langle{\hat{J}_{a_{x^{\prime},y^{\prime}}}}^{2}\rangle}.$ (54)
Similar to Eq. (34), we have
$\langle\hat{J}_{a_{x^{\prime}}}^{2}\rangle=\frac{1}{|\langle\hat{\mathbf{J}}_{a}\rangle|^{2}}\Big{[}\langle\hat{J}_{a_{x}}^{2}\rangle\langle\hat{J}_{a_{z}}\rangle^{2}+\langle\hat{J}_{a_{z}}^{2}\rangle\langle\hat{J}_{a_{x}}\rangle^{2}$
$-\langle\hat{J}_{a_{x}}\hat{J}_{a_{z}}+\hat{J}_{a_{z}}\hat{J}_{a_{x}}\rangle\langle\hat{J}_{a_{x}}\rangle\langle\hat{J}_{a_{z}}\rangle\Big{]}.$ (55)
We find that
$\langle\hat{J}_{a_{x}}^{2}\rangle=tr(\hat{\rho}_{a}\hat{J}_{a_{x}}^{2})=\frac{1}{4},$ (56)
$\langle\hat{J}_{a_{y}}^{2}\rangle=tr(\hat{\rho}_{a}\hat{J}_{a_{y}}^{2})=\frac{1}{4},$ (57)
$\langle\hat{J}_{a_{z}}^{2}\rangle=tr(\hat{\rho}_{a}\hat{J}_{a_{z}}^{2})=\frac{1}{4},$ (58)
$\langle\hat{J}_{a_{x}}\hat{J}_{a_{z}}+\hat{J}_{a_{z}}\hat{J}_{a_{x}}\rangle=tr\{\hat{\rho}_{a}(\hat{J}_{a_{x}}\hat{J}_{a_{z}}+\hat{J}_{a_{z}}\hat{J}_{a_{x}})\}$
$=0,$ (59)
Therefore, using Eqs. (45) to (I.2) and (54) to (59), we get
$\Delta J_{a_{x^{\prime}}}$ $=$ $\frac{1}{2},$ (60)
$\Delta J_{a_{y^{\prime}}}$ $=$ $\frac{1}{2}.$ (61)
Similarly, it can be shown that for the atom $b$ the uncertainties in $\hat{J}_{b_{x^{\prime}}}$ and $\hat{J}_{b_{y^{\prime}}}$ are
$\Delta J_{b_{x^{\prime}}}$ $=$ $\frac{1}{2},$ (62)
$\Delta J_{b_{y^{\prime}}}$ $=$ $\frac{1}{2}.$ (63)
Now, the von Neumann entropy of the composite state in Eq. (12) is zero as it is a pure state. Therefore,
$S_{a,b}=0.$ (64)
But, the von Neumann entropies of the atoms $a$ and $b$, which are entangled are
$S_{a}=S_{b}=-c_{1}^{2}log_{2}c_{1}^{2}-c_{2}^{2}log_{2}c_{2}^{2}.$ (65)
We now plot the uncertainties in Eqs. (42) and (43) and the von Neumann entropy in Eq. (65) with respect to the constant $c_{1}$ in FIG. 1.
Figure 1: Plot of the uncertainties $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$ and the von Neumann entropies $S_{a}$ and $S_{b}$ of the atoms $a$ and $b$ with respect to the constant $c_{1}$. Here we have taken $c_{3}=c_{4}=c_{5}=c_{6}=\frac{1}{\sqrt{2}}$. $c_{1}$ has been plotted along horizontal axis and $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$ and $S_{a}$ and $S_{b}$ along the vertical axis. The thin line, thick line and the dotted line represent $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$ and $S_{a}=S_{b}$ respectively.
We observe that though the von Neumann entropies $S_{a}$ and $S_{b}$ are always greater than 00, that is that of the composite state, in the whole range $0.001\leq c_{1}\leq 0.999$, as represented by the dotted line in FIG. 1, but both the uncertainties $\Delta J_{x^{\prime}}$ and $\Delta J_{y^{\prime}}$ are greater than $\frac{1}{2}$ in the range $0.001\leq c_{1}\leq 0.258$ and $0.966\leq c_{1}\leq 0.999$. The above ranges of $c_{1}$ correspond to the ranges of $c_{2}$ as $0.999\geq c_{2}\geq 0.966$ and $0.258\geq c_{2}\geq 0.001$. Therefore, both the uncertainties of the composite state in Eq. (12) are greater than those of the individual entangled atoms which have been represented in Eqs. (60) to (63). This is in contradiction with the prevailing consensus that more is the entropy, more is the uncertainty and more is the disorder.
Therefore, we see from FIG. 1 that there is violation of the entropic inequalities Horodecki1 , as in this case we have
$S_{a,b}<S_{a},$ (66)
$S_{a,b}<S_{b},$ (67)
but
$\Delta J_{x^{\prime}_{a,b}}>\Delta J_{x^{\prime}_{a}},$ (68)
$\Delta J_{x^{\prime}_{a,b}}>\Delta J_{x^{\prime}_{b}},$ (69)
and
$\Delta J_{y^{\prime}_{a,b}}>\Delta J_{y^{\prime}_{a}},$ (70)
$\Delta J_{y^{\prime}_{a,b}}>\Delta J_{y^{\prime}_{b}},$ (71)
in the ranges
$0.001\leq c_{1}\leq 0.258~{}~{}and~{}~{}0.966\leq c_{1}\leq 0.999,$
$0.999\geq c_{2}\geq 0.966~{}~{}and~{}~{}0.258\geq c_{2}\geq 0.001$
We have obtained this feature for the general state of two entangled two-level atoms.
We also observe from FIG. 1 that when the von Neumann entropies $S_{a}=S_{b}$ of the individual atoms attain their peak values, the uncertainty $\Delta J_{x^{\prime}}$ of the composite state attains its peak value while the uncertainty $\Delta J_{y^{\prime}}$ attains $100\%$ squeezing.
I.3 IV GENERALIZED RELATION BETWEEN VON NEUMANN ENTROPY AND SQUEEZING PARAMETERS IN TWO-LEVEL ATOMS
We establish generalized operational measures of the von Neumann entropies of the entangled atoms. We do this by establishing generalized relations between the von Neumann entropies of the entangled atoms and spin squeezing parameter introduced by Kitagawa et. al Kitagawa ; Jian and spectroscopic squeezing parameter used in the context of Ramsey spectroscopy, introduced by Wineland et. al Wineland2 ; Jian .
Using Eqs. (9) and (43), we get
$c_{1}^{4}-c_{1}^{2}=\Delta J_{y^{\prime}}^{2}-\Delta J_{y^{\prime}}^{4}-\frac{1}{4}$ (72)
This is a quadratic equation in $c_{1}^{2}$ whose roots are
$c_{1}^{2}=\frac{1}{2}\pm\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}$ (73)
Therefore, using Eqs. (9) and (73), we get
$c_{2}^{2}=\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}$ (74)
So, using Eqs. (65), (73) and (74), we get
$S_{a}$ $=$ $S_{b}=-\Big{[}\frac{1}{2}\pm\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}\Big{]}\times$ (75)
$log_{2}\Big{[}\frac{1}{2}\pm\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}\Big{]}-$
$\Big{[}\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}\Big{]}\times$
$log_{2}\Big{[}\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}\Big{]}$
or,
$S_{a}$ $=$ $S_{b}$
$=$ $-\Big{[}\frac{1}{2}\pm\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}\sqrt{\frac{j}{2}}\sqrt{1-\frac{2}{j}\Delta J_{y^{\prime}}^{2}\frac{j}{2}}\Big{]}\times$
$log_{2}\Big{[}\frac{1}{2}\pm\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}\sqrt{\frac{j}{2}}\sqrt{1-\frac{2}{j}\Delta J_{y^{\prime}}^{2}\frac{j}{2}}\Big{]}$
-- $\Big{[}\frac{1}{2}\mp\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}\sqrt{\frac{j}{2}}\sqrt{1-\frac{2}{j}\Delta J_{y^{\prime}}^{2}\frac{j}{2}}\Big{]}\times$
$log_{2}\Big{[}\frac{1}{2}\mp\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}\sqrt{\frac{j}{2}}\sqrt{1-\frac{2}{j}\Delta J_{y^{\prime}}^{2}\frac{j}{2}}\Big{]}.$
Here $j=\frac{N}{2}$, where $N$ is the number of atoms. In this case we have $N=2$ and hence $j=1$. Now,
$\sqrt{\frac{2}{j}}\Delta J_{\overrightarrow{n}_{\perp}}=\xi_{s}$ (77)
is the spin squeezing parameter introduced by Kitagawa et. al Kitagawa ; Jian , where $\overrightarrow{n}_{\perp}$ is in the plane perpendicular to the mean spin vector $\langle\hat{\mathbf{J}}\rangle$.
In this case the $\overrightarrow{n}_{\perp}$ are the $x^{\prime}$ and $y^{\prime}$ directions. Now we have observed from FIG. 1 that we have squeezing in the $y^{\prime}$-quadrature. So, we take $\Delta J_{\overrightarrow{n}_{\perp}}=\Delta J_{y^{\prime}}$. Therefore, in this case
$\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}=\xi_{s_{y}},$ (78)
and hence we get with $j=1$,
$S_{a}$ $=$ $S_{b}=-\Big{[}\frac{1}{2}\pm\frac{{\xi_{s_{y}}}}{\sqrt{2}}\sqrt{1-\frac{\xi_{s_{y}}^{2}}{2}}\Big{]}\times$
$log_{2}\Big{[}\frac{1}{2}\pm\frac{{\xi_{s_{y}}}}{\sqrt{2}}\sqrt{1-\frac{\xi_{s_{y}}^{2}}{2}}\Big{]}$
-- $\Big{[}\frac{1}{2}\mp\frac{{\xi_{s_{y}}}}{\sqrt{2}}\sqrt{1-\frac{\xi_{s_{y}}^{2}}{2}}\Big{]}\times$
$log_{2}\Big{[}\frac{1}{2}\mp\frac{{\xi_{s_{y}}}}{\sqrt{2}}\sqrt{1-\frac{\xi_{s_{y}}^{2}}{2}}\Big{]}.$
The spectroscopic squeezing parameter $\xi_{R}$, introduced by Wineland et. al Wineland2 ; Jian is
$\xi_{R}=\sqrt{N}\frac{\Delta J_{\overrightarrow{n}_{\perp}}}{|\langle\hat{\mathbf{J}}\rangle|}.$ (80)
In this case since $N=2$, and we have squeezing in the $y^{\prime}$-quadrature, we have
$\xi_{R_{y}}=\sqrt{2}\frac{\Delta J_{y^{\prime}}}{|\langle\hat{\mathbf{J}}\rangle|}.$ (81)
We write Eq. (75) as
$S_{a}$ $=$ $S_{b}$
$=$ $-\Big{[}\frac{1}{2}\pm\xi_{R_{y}}\frac{|\langle\hat{\mathbf{J}}\rangle|}{\sqrt{2}}\sqrt{1-\xi_{R_{y}}^{2}\frac{|\langle\hat{\mathbf{J}}\rangle|^{2}}{2}}\Big{]}\times$
$log_{2}\Big{[}\frac{1}{2}\pm\xi_{R_{y}}\frac{|\langle\hat{\mathbf{J}}\rangle|}{\sqrt{2}}\sqrt{1-\xi_{R_{y}}^{2}\frac{|\langle\hat{\mathbf{J}}\rangle|^{2}}{2}}\Big{]}$
-- $\Big{[}\frac{1}{2}\mp\xi_{R_{y}}\frac{|\langle\hat{\mathbf{J}}\rangle|}{\sqrt{2}}\sqrt{1-\xi_{R_{y}}^{2}\frac{|\langle\hat{\mathbf{J}}\rangle|^{2}}{2}}\Big{]}\times$
$log_{2}\Big{[}\frac{1}{2}\mp\xi_{R_{y}}\frac{|\langle\hat{\mathbf{J}}\rangle|}{\sqrt{2}}\sqrt{1-\xi_{R_{y}}^{2}\frac{|\langle\hat{\mathbf{J}}\rangle|^{2}}{2}}\Big{]}.$
The Eqs. (LABEL:3.57) and (LABEL:3.60) give generalized operational measures of von Neumann entropies of the entangled atoms.
Figure 2: Plot of the spin squeezing parameter $\xi_{s_{y}}^{2}$, spectroscopic squeezing parameter $\xi_{R_{y}}^{2}$ and the von Neumann entropies $S_{a}$ and $S_{b}$ of the entangled parts with respect to the constant $c_{1}$. Here we have taken $c_{3}=c_{4}=c_{5}=c_{6}=\frac{1}{\sqrt{2}}$. $c_{1}$ has been plotted along the horizontal axis and $\xi_{s_{y}}^{2}$, $\xi_{R_{y}}^{2}$ and $S_{a}=S_{b}$ along the vertical axis. The thick line, dashed line and the dotted line represent $\xi_{s_{y}}^{2}$, $\xi_{R_{y}}^{2}$ and $S_{a}=S_{b}$ respectively.
We see from FIG. 2 that as the spin squeezing and spectroscopic squeezing of the composite state increases the von Neumann entropies of the entangled parts increase and when there is $100\%$ spin squeezing of the composite state the von Neumann entropies of the entangled parts attain their peak values. Therefore, more is the squeezing of the composite state more is the von Neumann entropy of the entangled parts. Then as the squeezing of the composite state decreases the von Neumann entropy of the entangled parts decrease. This feature, as we see, is for the general state of two two-level atoms given in Eq. (12).
I.4 V CONCLUSION
We have calculated analytically the uncertainties for a general state of two entangled two-level atoms both for the composite state and the entangled parts. We observe that the von Neumann entropies of the entangled parts are always greater than that of the composite state, but for certain range of the superposing constants of the quantum state, the uncertainties of the entangled parts are less than that of the composite state. This is not in accordance with the prevailing consensus that more is the entropy, more is the uncertainty. This feature, as we see, is for the general state of two two-level atoms.
We have also established analytically the generalized relationship between the von Neumann entropies of the entangled parts and the uncertainties, spin squeezing and spectroscopic squeezing parameters, used in the context of Ramsey spectroscopy, of the composite state. This provides a generalized operational measure of the von Neumann entropies of the entangled atoms.